let z1=8(cos5pi/4 + isin5pi/4) and z2=2(cos5pi/12 + isin5pi/12)
Find each of the following in the form a+bi
a. z1/z2
b. (z2)^6

Answer by stanbon(57347)

let z1=8(cos5pi/4 + isin5pi/4) and z2=2(cos5pi/12 + isin5pi/12)
Find each of the following in the form a+bi
a. z1/z2
= (4)(cos[(15pi-5pi)/12 + isin[same])
= 4(cos[(5/6)pi] + isin[same])
= -3.464 + 2i
======================
b. (z2)^6
= (2^6)(cos[6[5pi/12]] + isin[same])
= 64(cos[(5/2)pi] + isin[same])
= 0 +64i
=================
Cheers,
Stan H.